OpenStudy (rock_mit182):

Mary runs 3 miles in 25 minutes. At the same rate, how many miles would she run in 20 minutes?

OpenStudy (mathmale):

Hallo! Wie geht's? suggest you find Mary's speed (rate) first. It is v = (3 mi) / (25 min). Remember, distance = rate times time. Thus, having calculated the rate, just multiply it by 20 minutes, cancelling out "minutes" as appropriate.

OpenStudy (rock_mit182):



OpenStudy (rock_mit182):

is it right ?

OpenStudy (rock_mit182):

PLease i need to answer the last one right ! and the time is running, actually

OpenStudy (rock_mit182):

I Just need to verify, am I right ?

OpenStudy (rock_mit182):

x = 2,4 miles ?

OpenStudy (doc.brown):

You can also think of it as a conversion. You are converting 20min to miles using a rate. \[20min\times\frac{3mi}{25min}\]\[20\cancel{min}\times\frac{3mi}{25\cancel{min}}\]\[\frac{20\times 3mi}{25}\]\[=2.4mi\]

OpenStudy (rock_mit182):

OK thanks bro I was just checking

OpenStudy (mathmale):

Es tut mir Leid, dass ich so langsam bin. Du hast jawohl Recht.

OpenStudy (rock_mit182):

Nein, ist ok bruder, dankeschön !

OpenStudy (rock_mit182):

@mathmale I appreciate your help, thumbs up

OpenStudy (mathmale):

Bis bald!